Point charges Q_1 = + 3 10^6 C and Q_2 = -9 10^-6d
Answer #1

a) 16 x=-5.6 V = 0.0b) The point is closer to Q1 as you can see in thr graph.

The electric field is given by the rate of change of the electric potential respect to the position. The electric field is zero if and only if the electric potential is constant. In the point the electric potential is clearely not constant, therefore the electric field at that point is not zero

c) The total electric potential at any point in the x axis is

V(x)=k_e\frac{q_1}{|x+11|}-k_e\frac{q_2}{|x-11|}

for the potential zero

0=k_e\frac{q_1}{|x+11|}-k_e\frac{q_2}{|x-11|}

k_e\frac{q_2}{|x-11|}=k_e\frac{q_1}{|x+11|}

{q_1}{|x-11|}={q_2}{|x+11|}

Her we have to solve two equations

q_1(x-11)=q_2(x+11)

q_1x-11q_1=q_2x+11q_2

-11q_1-11q_2=q_2x-q_1x

-11(q_1+q_2)=(q_2-q_1)x

x=-11\frac{q_1+q_2}{(q_2-q_1)}

x=-11\frac{3+9}{9-3}

x=22m

but that point is not between the charges

the second equation is

q_1(11-x)=q_2(x+11)

11q_1-q_1x=q_2x+11q_2

11q_1-11q_2=q_2x+q_1x

11(q_1-q_2)=(q_2+q_1)x

x=11\frac{q_1-q_2}{(q_2+q_1)}

x=11\frac{3-9}{9+3}

x=-5.5m

That is the point

d) The potential energy like is not defined. But in electrostatic is common to reffere to this potential energy at one point the difference of potential energy from the infinite to the point. That potential energy is given by

U=qVfor point charges.

Since the elecetric potential is zero at that point

U=q(0)

U=0J

The kinetic energy is

K=\frac{1}{2}mv^2

K=\frac{1}{2}m(0)^2

K=0J

The charge is pushed by the positive charge in +x direction and pulled by the negative charge in +x direction. The charge will moves to +x direction. Note that althought the potentail energy is zero, the charge is not in an equillibrium point.

e) At the origin the electrip potential is

V(x)=(8.99\times 10^{9}N\cdot m^2/C^2)\frac{(3\times 10^{-6}C)}{|(0)+11m|}-(8.99\times 10^{9}N\cdot m^2/C^2)\frac{(9\times 10^{-6}C)}{|(0)-11m|}V(x)=-4900V

Then the change in the potential energy is

\Delta U=(2\times 10^{-3}C)(-4900V-0V)

\Delta U=-9.8J

And by the conservation of the energy

\Delta K=-\Delta U

(K_2-K_1)=-(-9.8J)

(K_2-(0J))=-(-9.8J)

K_2=9.8J